A company has long-term borrowings of K3,000,000 on which it pays interest at 6% per year. The equity capital consists of 500,000 of K10 each. The share premium account stands at K2,500,000 and retained earnings amount to K2,000,000. Its shares are feely traded on the stock exchange and buyers’ and sellers’ quotes have remained steady at K21 and K22 respectively. Some sales of shares have been recorded at K21.20. The company has paid a stable dividend of 10% for the last 10 years. The market price of debt is 8%. The company is expanding rapidly and requires an additional K4,000,000 long-term capital.



Required:

Compute the present weighted cost of capital.                                              



1
Expert's answer
2021-04-12T13:10:21-0400

WACC calculation.


"WACC=(Ke\times \frac{E}{V}) + (Kd\times\frac{D}{V})"


Where; V= Total debt(D) + Total equity (E), Ke is cost of equity and Kd is cost of debt


Total debt=3,000,000


"Kd=After tax cost=0.7\times6=4.2"


Total equity=K 5,000,000+K 2,500,000+K 2,000,000=K9,500,000


Ke =10%


V=9,500,000+3,000,000=12,500,000


"WACC=(10\times \frac{9,500,000}{12,500,000}) + (4.2\times\frac{3,000,000}{12,500,000})"

WACC==7.6+1.008=8.608%